### v1

@@ I Introduction
To facilitate the study, we collected 2 million Python codes within six years (January 1st, 2018, to December 31st, 2023). The contaminated data is a form of the codes whose timestamps precede models’ cut-off dates. The cleansed data are derived from implementing countermeasures ➊ $\sim$ ➌. Furthermore, we use three code similarities (i.e., Jaccard distance, Levenshtein distance, and CodeBleu as used in related work [26]) to quantify the overlap degree between contaminated/cleansed data, showing that the overlap is at a relatively minimal level (see Section IV-A). In addition, we compute the code similarities between contaminated and cleansed data to make sure that the overlap between them are under a relatively small level. We study three research questions (RQs) accordingly:
@@ II-A Experiment Design
To facilitate the study, we construct several groups of code (code groups) into contaminated and cleansed data. Comparing the CLM’s performance on the contaminated and cleansed data can infer the extent to which the CLMs rely on memorization versus their genuine ability, and thus act as a hint for the degree of data contamination during the evaluation. Fig. 1 visualizes an overview design for RQ1-RQ3. The cleansed data (highlighted in yellow) could be collected after model release (RQ1), curated manually (RQ2), or mutated from contaminated data (RQ3). These three RQs study how CLMs’ performance changes over contaminated/cleased data. Finally, RQ4 explores whether there are metrics that could distinguish them.
@@ II-A1 Experiment Design for RQ1
For clarity, we designate code within a specific year with corresponding labels, such as Code-2018 for snippets that fall in 2018, and similarly for other annual collections. Note that the pivot time of contaminated/recent data varies from CLMs (See Section III-A).
After the temporal splitting, we apply CLMs to these code groups, allowing us to systematically investigate how CLMs’ performance changes chronologically, thereby implying the degree of data contamination within these CLMs.
@@ II-A2 Experiment Design for RQ2
The release time of HumanEval and CoderEval are Jul 8, 2021 [32] and Jan 28, 2023 [33], respectively, according to the commit date shown in Github. We then compare these commit dates with models’ cut-off date, as evidence of whether these datasets are contaminated data with regard to a CLM or not.
@@ II-A3 Experiment Design for RQ3
As proposed by existing works [9], a more applicable way to alleviate the threat of data contamination in evaluating CLM-based approaches is to refactor the contaminated code. If CLMs maintain performance on refactored code, it would imply that these CLMs truly grasp the coding tasks rather than relying on rote memorization. In contrast, if there is a substantial fluctuation in CLMs’ performance between contaminated and refactored data, it indicates that CLMs are likely threatened by data contamination via memorizing specific code patterns or formats in the training data. Conversely, if a refactoring operator can induce significant variability (primarily decreases) in CLMs’ performance, it may mitigate data contamination to which the CLM is subject.
@@
In particular, on the one hand, syntactic operators check whether CLMs rely on rote memorization of pattern recognition. The refactoring preserves the semantic purpose but alters its syntactic structure [34, 35]. If CLMs truly understand the semantics of the code, their performance should not significantly decline when faced with semantically equivalent but syntactically varied code snippets. On the other hand, the semantic operators ensure CMLs understand code semantics rather than keyword/string matching. Replacing with synonyms or adding more semantics should not significantly affect model performance.
@@ II-E2 Semantic Refactoring Operators
  • Performance Measurement Decoration (Deco). Since measuring the performance of a function, e.g., memory usage and execution time do not change the function’s behavior, we also refactor the Python functions by adding performance measuring decorators to functions: @timing, which measures the total execution time of the decorated function, and @measure_memory_usage, which measures the memory usage of the function.
@@ III-B1 Data Collection from the Stack v2
We take the Stack v2 [26] dataset as the code corpus for its representativeness and comprehensiveness. It contains over 6.4TB [^2] We use the deduplicated version of the stack which contains over 2.9TB data to avoid code duplication. of permissively licensed source code files covering 358 programming languages, from January 1st, 2015 to March 31st, 2022. Typically, we take Python as our subject programming language for its popularity.
@@ III-D Curated Code Groups Preparation (RQ2)
HumanEval [17] dataset contains a set of Python functions, so we construct the Code-HumE code group directly with all functions in HumanEval. CoderEval [20] dataset contains a set of Python files, so we first extract functions from all files in the same way as in Section III-B3 and then construct the Code-CodE code group.
@@ III-E Refactored Code Groups Preparation (RQ3)
The refactoring operators are expected to apply to contaminated code. According to Table I, the earliest cut-off date is September 2021, so we chose Code-2021 to apply the refactoring. In particular, we apply code refactoring operators introduced in Section II-E to Code-2021. Each operator is applied to every function in Code-2021 where applicable. If an operator is not applicable (e.g.., when attempting to flip if-else branches in code that does not contain an if-statement) to a certain code function under refactoring, we skip refactoring that function.
If there are multiple ways to apply a certain operator, we randomly select one way to apply it. For example, if one function has multiple if-conditions, we randomly pick one if-condition and apply the IFF (if-condition flipping) operator to it. The rationale behind employing a single operator at most once without repetition is that we want to analyze the impact of the most granular refactoring operator. Besides, by applying at most one type of operator to one function, we aim to isolate the influence of an individual operator rather than the compounded effects resulting from multiple operators. We use Code-IFF, Code-Loop, Code-Renm, Code-Param, and Code-Deco to refer to the code group constructed by applying the five refactoring operations in Section II-E, respectively.
@@ IV-A Evaluation on the Assumption of RQ1-RQ3
In general, the contaminated/cleansed code groups share similar scores in all metrics. The mean perplexity of the code groups ranges from 3.37 (Code-2023) to 5.17 (Code-CodE). The largest score difference is observed in ppl_lower, ranging from 4.06 (Code-2023) to 6.41 (Code-CodE). In addition, the Code-2023, Code-CodE, and the five refactored code groups are unseen for StarCoder whose cut-off date is March 2022, yet there is no apparent difference in these cleansed code groups compared with that on contaminated code groups.
@@
[style=MyFrame] Finding 8: The existing six MIA-related metrics such as perplexity can hardly distinguish the contaminated/cleansed data, so they may be hardly used for quantifying data contamination.
Surprisingly, though the recent code, i.e., Code-2023, is later than StarCoder’s cut-off date (March 2022), all six metrics show absolute drops on it compared with other code groups. It means that Code-2023 fits CLMs’ training distribution better than other code groups, so the model found it more natural (lower scores). This observation may partially explain the counter-intuitiveness we observed in RQ1. CLMs find Code-2023 is more natural than earlier code groups and is closer to the training distribution; their performance will naturally be better.

### v2

@@ I Introduction
To facilitate the study, we collected 2 million Python codes within six years (January 1st, 2018, to December 31st, 2023). The contaminated data is a form of the codes whose timestamps precede models’ cut-off dates. The cleansed data are derived from implementing countermeasures ➊ $\sim$ ➌. Furthermore, we use three code similarities (e.g., Jaccard distance, Levenshtein distance as used in related work [26]) to quantify the overlap degree between contaminated/cleansed data, showing that the overlap is at a relatively minimal level (Section IV-A). In addition, we compute the code similarities between contaminated and cleansed data to make sure that the overlap between them is under a relatively small level. We study three research questions (RQs):
@@ II-A Experiment Design
To facilitate the study, we construct several groups of code (code groups) into contaminated and cleansed data. Comparing the CLM’s performance on the contaminated and cleansed data can infer the extent to which the CLMs rely on memorization versus their genuine ability and thus act as a hint for the degree of data contamination during the evaluation. Fig. 1 visualizes an overview design for RQ1-RQ3. The cleansed data (highlighted in yellow) could be collected after model release (RQ1), curated manually (RQ2), or mutated from contaminated data (RQ3). These three RQs study how CLMs’ performance changes over contaminated/cleansed data. Finally, RQ4 explores whether there are metrics that could distinguish them.
@@ II-A1 Experiment Design for RQ1
For clarity, we designate code within a specific year with corresponding labels, such as Code-2018 for snippets that fall in 2018 and similarly for other annual collections. Note that the pivot time of contaminated/recent data varies from CLMs (See Section III-A).
After the temporal splitting, we apply CLMs to these code groups, allowing us to systematically investigate how CLMs’ performance changes chronologically, implying the degree of data contamination within them.
@@ II-A2 Experiment Design for RQ2
The release time of HumanEval and CoderEval are Jul 8, 2021 [32] and Jan 28, 2023 [33], respectively, according to the commit date shown in Github. We then compare these commit dates with models’ cut-off dates to determine whether these datasets contain contaminated data regarding a CLM.
@@ II-A3 Experiment Design for RQ3
As proposed by existing works [9], a more applicable way to alleviate the threat of data contamination in evaluating CLM-based approaches is to refactor the contaminated code. If CLMs maintain performance on refactored code, this would imply that they truly grasp the coding tasks rather than relying on rote memorization. In contrast, if there is a substantial fluctuation in CLMs’ performance between contaminated and refactored data, it indicates that CLMs are likely threatened by data contamination via memorizing specific code patterns or formats in the training data. Conversely, if a refactoring operator can induce significant variability (primarily decreases) in CLMs’ performance, it may mitigate the CLM’s exposure to data contamination.
@@
On the one hand, syntactic operators check whether CLMs rely on rote memorization of pattern recognition. The refactoring preserves the semantic purpose but alters its syntactic structure [34, 35]. If CLMs truly understand the semantics of the code, their performance should not significantly decline when faced with semantically equivalent but syntactically varied code snippets. On the other hand, the semantic operators ensure CMLs understand code semantics rather than keyword/string matching. Replacing with synonyms or adding more semantics should not significantly affect model performance.
@@ II-E2 Semantic Refactoring Operators
  • Performance Measurement Decoration (Deco). Since measuring the performance of a function, e.g., memory usage and execution time do not change the function’s behavior, we also add performance measuring decorators to functions: @timing, which measures the total execution time of the decorated function, and @measure_memory_usage, which measures the memory usage of the function.
@@ III-B1 Data Collection from the Stack v2
We take the Stack v2 [26] dataset as the code corpus for its representativeness and comprehensiveness. We use the deduplicated version of the stack, which contains over 2.9TB data of permissively licensed source code files covering 358 programming languages, from January 1st, 2015 to March 31st, 2022. Typically, we take Python as our subject programming language for its popularity.
@@ III-D Curated Code Groups Preparation (RQ2)
HumanEval [17] dataset contains a set of Python functions, so we construct the Code-HumE code group directly with all functions in HumanEval. CoderEval [20] dataset contains a set of Python files, so we extract functions from all files like in Section III-B3 and then construct the Code-CodE code group.
@@ III-E Refactored Code Groups Preparation (RQ3)
The refactoring operators are expected to apply to contaminated code. According to Table I, the earliest cut-off date is September 2021, so we chose Code-2021 to apply the refactoring. In particular, we apply the operators introduced in Section II-E to Code-2021. Each operator is applied to every function in Code-2021 where applicable. If an operator is not applicable (e.g.., when attempting to flip if-else branches in code that does not contain an if-statement) to a certain code function under refactoring, we skip refactoring that function.
If there are multiple ways to apply a particular operator, we randomly select one way to use it. For example, if one function has multiple if-conditions, we randomly pick one if-condition and apply the IFF (if-condition flipping) operator. The rationale behind employing a single operator at most once without repetition is that we want to analyze the impact of the most granular refactoring operator. Besides, by applying at most one type of operator to one function, we aim to isolate the influence of an individual operator rather than the compounded effects resulting from multiple operators. We use Code-IFF, Code-Loop, Code-Renm, Code-Param, and Code-Deco to refer to the code group constructed by applying the five refactoring operations in Section II-E, respectively.
@@ IV-A Evaluation on the Assumption of RQ1-RQ3
The contaminated/cleansed code groups generally share similar scores in all metrics. The mean perplexity of the code groups ranges from 3.37 (Code-2023) to 5.17 (Code-CodE). The largest score difference is observed in ppl_lower, ranging from 4.06 (Code-2023) to 6.41 (Code-CodE). In addition, the Code-2023, Code-CodE, and the five refactored code groups are unseen for StarCoder whose cut-off date is March 2022, yet there is no apparent difference in these cleansed code groups compared with that on contaminated code groups.
@@
[style=MyFrame] Finding 8: The existing six MIA-related metrics can hardly distinguish the contaminated/cleansed data, so they may be hardly used for quantifying data contamination.
Surprisingly, though the recent code, i.e., Code-2023, is later than StarCoder’s cut-off date (March 2022), all six metrics show absolute drops on it compared with other code groups. Code-2023 fits CLMs’ training distribution better than other code groups, so the model found it more natural (lower scores). This observation may partially explain the counterintuitiveness we observed in RQ1. CLMs find Code-2023 more natural than earlier code groups and closer to the training distribution; their performance will naturally be better.